# Characterizing the Sensitivity to Individual Bit Flips in Client-Side Operations of the CKKS Scheme

### v1

@@ III-D2 Bit‐Flip in the Ciphertext
For a perturbed ciphertext $ct^{\prime}=(c_{0}+e_{j}\,X^{i-1},\;c_{1})$, the behavior mirrors the previous plaintext case: the error remains localized in coefficient $i$ and results in a decoding error proportional to $e_{j}$.
@@
For a perturbed ciphertext $ct^{\prime}=(c_{0},\;c_{1}+e_{j}\,X^{i-1})$, decryption proceeds as:
$\begin{split}\boldsymbol{m}^{\prime\prime}&=[\,c_{0}+(c_{1}+e_{j}\,X^{i-1})\,s\,]_{Q}\\ &=[\,(c_{0}+c_{1}\,s)+(e_{j}\,X^{i-1}\,s)\,]_{Q}\\ &=[\,(\boldsymbol{m}+e_{1}+e_{2}\,s)+(e_{j}\,X^{i-1}\,s)\,]_{Q}\end{split}$
where $e_{j}X^{i-1}\,s$ involves polynomial multiplication. Thus, the error spreads across all $N$ coefficients. Since the secret key $s$ has small coefficients (much less than $q_{0}$), each coefficient in the result incurs an error of magnitude proportional to $e_{j}$, but this error is now distributed. The maximum magnitude is bounded by $|e_{j}|\cdot\|s\|_{\infty}$.
III-D3 Bit‐Flip in the Plaintext Before Decoding
If a bit-flip is introduced in coefficient $i$ of a plaintext $\boldsymbol{m}$, after decryption but before decoding, the effect is analogous to the case analyzed in Section III-C. This follows directly since the plaintext being decoded is equivalent to:
$\boldsymbol{m}^{\prime}=\boldsymbol{m}+e_{j}\,X^{i-1}$ (15)
@@ III-E Theoretical Estimation of Error in RNS
$r_{k}=p\bmod q_{k}\quad\text{for }k=0,\dots,L$
When applied to all coefficients of a polynomial (from an encoded message or ciphertext), this yields $L$ residue polynomials, commonly referred to as limbs. Reconstruction via the Chinese Remainder Theorem (CRT), as shown in Mohan [25, Chapter 5], is:
@@
where
$Q_{k}=Q/q_{k}$
and $\frac{1}{Q_{k}}$ are the multiplicative inverse of $Q_{k}$.
A bit-flip on the $j$ bit affecting the $k$-th residue changes it to $r_{k}^{\prime}=r_{k}+e_{k,j}$, which introduces an error into the reconstructed value:
$p^{\prime}=p+e_{k,j}\,Q_{k}\,\left[\left(\frac{1}{Q_{k}}\right)\bmod q_{k}\right]\pmod{Q}$ (17)
This error can be as large as $Q_{k}$, potentially spanning hundreds or thousands of bits, even though the bit-flip affected only one bit in a single residue.
@@ III-F Theoretical Estimation of Error in NTT
$\psi^{2}\equiv\xi,\quad\psi^{N}\equiv-1\pmod{q},$
where $\xi$ is a primitive $N$-th root of unity. The NTT matrix is then defined as:
We denote its inverse as $W^{-1}_{NTT}$. When a single bit-flip occurs at bit position $j$ within coefficient $i$, the induced perturbation after NTT transformation becomes:
$\mathbf{m}^{\prime}-\mathbf{m}=W_{NTT}\,\mathbf{e}_{i,j}.$
Therefore, the $L_{2}$ norm of the error is:
[…]
Figure 6 shows the values of $L_{1}(i,j)$ for $N=4$ and 64-bit coefficients in the NTT case. The horizontal axis indicates the bit-flip position $i\cdot 64+j$, and the vertical axis shows $L_{2}(i,j)$ on a logarithmic scale.
Figure 6: $L_{2}$ norm of the error after a single bit-flip in the plaintext during encoding. The horizontal axis shows the modified bit position $i*64+j$; the vertical axis shows the $L_{2}(i,j)$ norm on a logarithmic scale. Ring dimension: $N=4$.
As the Fig. 6 shows, even the flips in the least significant bits cause the norm to exceed four orders of magnitude—producing errors large enough to make correct decoding practically impossible, as will be further illustrated in Section V.
III-G Summary
The following points summarize the theoretical impact of a single bit-flip at different stages of the CKKS encryption pipeline, both in its vanilla and optimized implementations:
  • Bit-flip in the plaintext before encryption: the error remains localized in a single coefficient where the bit-flip occurred, with a magnitude proportional to $e_{j}=2^{j}$, depending on the flipped bit’s position $j$.
  • Bit-flip in the ciphertext ($c_{0}$): identical to the plaintext case—the error remains localized after decryption, with magnitude again proportional to the bit position $j$.
  • Bit-flip in the ciphertext ($c_{1}$): the error is multiplied by the secret key $s$, becoming distributed across all $N$ coefficients; its maximum magnitude is $\mathcal{O}(e_{j}\|s\|_{\infty})$.
  • Bit-flip in the plaintext before decoding: behaves like a localized error in a single coefficient of the encoded polynomial–effectively equivalent to the encoding case, plus minor encryption noise.
  • RNS: a bit-flip in a single limb (residue) can result in a reconstruction error of order $\mathcal{O}(Q_{k})$, where $Q_{k}=Q/q_{k}$ is typically a large value.
[…]
IV Experimental Design
@@
This section describes the experimental design used to quantify the resilience of the CKKS homomorphic encryption scheme under single bit-flip fault injections.
IV-A Test Environment
[…]
  • HEAAN 1.0: a vanilla implementation without RNS or NTT, using arbitrary-precision coefficients managed by the NTL library (a C++ library for number theory and arbitrary-precision arithmetic).
  • OpenFHE 1.2.3: a modular implementation with native support for RNS and NTT, using 64-bit internal coefficients (uint64).
Although more recent versions of HEAAN include RNS and NTT optimizations, version 1.0 allows a cleaner comparison to a vanilla CKKS. Conversely, OpenFHE represents the optimized variant, due to its widespread use in the field of homomorphic encryption [26] and its active support for high-performance execution environments.
@@ IV-A Test Environment
Minor modifications were made to both libraries to gain full control over the selection of pseudo-random number generator (PRNG) seeds. This ensured strict reproducibility of experiments and allowed for precise isolation of the effects caused by each individual bit-flip.
In the case of HEAAN 1.0, we have these additional implementation details:
@@
  • The bit-length of each coefficient is not fixed in advance; it depends on internal combinations of $q_{0}$ and $\Delta$, and may vary depending on the PRNG seed used by the scheme.
@@
  • Due to NTL implementation, bits larger than $Q$ are ignored in internal operations, which limits the scope of bit-level analysis.
OpenFHE has the following implementation details:
  • Uses 64-bit integer coefficients (uint64_t) without arbitrary-precision arithmetic.
@@
  • The initial modulus $q_{0}$ can be explicitly configured by the user. This parameter determines the starting bit-level precision of the scheme. The subsequent moduli $q_{i}$ are automatically selected to match the bit length of $q_{0}$, ensuring uniform precision across levels.
@@
  • The total number of moduli depends on the required multiplicative depth of the computation.
  • Public APIs for accessing and modifying coefficients at each pipeline stage (i.e., encoding, encryption, etc.).
  • Enables injection of fixed-length bit-flips (up to 64 bits) at controlled and reproducible positions.
@@
  • Supports 128-bit arithmetic, which–though not used in this work–offers a pathway for future experiments.
IV-B Using OpenFHE as Vanilla CKKS implementation
OpenFHE is significantly more computationally efficient than HEAAN. However, it does not natively offer a mechanism to disable RNS and NTT optimizations. To address this, we devised a procedure to emulate a vanilla CKKS scheme within OpenFHE by explicitly bypassing both RNS and NTT optimizations, despite their built-in support in the library:
  • RNS: We enforced a single-limb representation, which is feasible for depth-zero operations (i.e., no ciphertext multiplications). Additionally, automatic rescaling was disabled and replaced with manual scaling to prevent the introduction of additional limbs.
  • NTT: Prior to injecting a bit-flip, the polynomial is transformed from evaluation space back to coefficient space via inverse NTT (INTT). We flip the desired bit and transform the polynomial back to the evaluation space using NTT. After this hook, the pipeline continue normally.
@@ IV-B Using OpenFHE as Vanilla CKKS implementation
  • Limitations of disabling NTT: Since the NTT operates modulo $q_{0}$, flipping bits beyond the bit-width of $q_{0}$ causes modular wraparound, which can alter the intended effect of the bit-flip and potentially lead to catastrophic decryption errors. To minimize distortion in the analysis, we set $q_{0}$ to the maximum value supported by the library–60 bits–ensuring that most bit-flips occur within the valid range and reducing unintended side effects from modular arithmetic.
IV-C Internal Error Detection in OpenFHE
OpenFHE includes an internal consistency check that estimates the standard deviation of the imaginary part of the decoded plaintext, designed to detect potential secret key leakage attacks. Flipping high-order bits at different stages of the homomorphic pipeline can result in large deviations, triggering this mechanism and aborting the decoding. To isolate the analysis of numerical resilience, we disable this mechanism.
This methodology enables controlled and reproducible exploration of both vanilla and optimized CKKS configurations, allowing for a rigorous assessment of the impact of single-bit faults at various stages of the homomorphic pipeline.
IV-D Vanilla CKKS over OpenFHE Validation
To ensure that OpenFHE’s vanilla mode faithfully replicates the behavior of HEAAN 1.0, we following these validation steps:
  1. 1.
    Configure OpenFHE in vanilla mode (not using RNS and NTT) and set $q$ and $\Delta$ to match HEAAN’s effective bit precision.
  2. 2.
    Restrict input values in HEAAN to real numbers, in line with OpenFHE’s plaintext constraints.
  3. 3.
    Compare $L_{2}$ norms and error statistics after identical bit-flips in both libraries.
  4. 4.
    Verify consistent sensitivity patterns across low- and high-order bit positions.
IV-E Input Vector Generation and PRNG Control
Since our analysis is not tied to any specific application, the input data consists of real-valued vectors with uniformly distributed random entries. To evaluate the sensitivity of CKKS, we define input ranges using power-of-two boundaries and sample values uniformly within these intervals.
A wider input range affects the signal-to-noise ratio by increasing the difference between the smallest and largest representable values.
Unless otherwise specified, input values are sampled as uniformly distributed real numbers within the interval 0 to 2560256start_ARG 0 end_ARG to start_ARG 256 end_ARG, and represented in double-precision floating-point format. The bounded input range allows us to test multiple $\Delta$, allowing for a more comprehensive analysis of their impact on encoding and rounding errors. By varying $\Delta$ across a wide range without exceeding coefficient bounds, we can better understand the trade-offs between precision and noise growth, as discussed in Section II-B.
To evaluate stability with respect to seed variation, each experiment (i.e., each individual bit-flip) is repeated 2500 times. Each repetition involves 100 different seeds for the scheme PRNG and 25 for the input generator. These seed combinations are fixed and shared across all experiments, ensuring that results could be compared under equivalent input conditions.
Although this number of repetitions may not suffice for exhaustive statistical inference, our tests show very low variability, confirming it is enough for the aims of this study.
IV-F Bit-Flip Injection Procedure
Error injection is performed by flipping a single bit in one encrypted coefficient during each execution. The general workflow for the experiments is as follows:
@@ IV-F Bit-Flip Injection Procedure
    (If required) Apply the inverse NTT (INTT) to bring the plaintext or ciphertext into coefficient space.
@@
    (If required) Reapply NTT to return to evaluation space.
@@
This process is performed once per bit and coefficient across both plaintext and ciphertext scenarios in separate experiments. The entire procedure was repeated for each seed combination.
In OpenFHE, we test configurations both with and without RNS and NTT by dynamically disabling these optimizations at runtime. This allows direct comparisons between the optimized and vanilla implementations, isolating the impact of each optimization technique.
IV-G Evaluation Metrics
Given the original decrypted vector $\mathbf{x}$ (without bit-flips) and the resulting vector $\mathbf{y}$ after error injection and decryption, we quantify the deviation introduced by each bit-flip by two complementary evaluation metrics:

### v3

@@ III-D2 Bit Flip in the Ciphertext
For a perturbed ciphertext $ct^{\prime}=(c_{0}+\boldsymbol{e}_{i-1,j},\;c_{1})$, the behavior mirrors the previous plaintext case: the error remains localized in coefficient $i$ and results in a decoding error proportional to $e_{j}$.
@@
For a perturbed ciphertext $ct^{\prime}=(c_{0},\;c_{1}+\boldsymbol{e}_{i-1,j})$, decryption proceeds as:
$\begin{split}\boldsymbol{m}^{\prime\prime}&=[\,c_{0}+(c_{1}+\boldsymbol{e}_{i-1,j})\,s\,]_{Q}\\ &=[\,(c_{0}+c_{1}\,s)+(\boldsymbol{e}_{i-1,j}\,s)\,]_{Q}\\ &=[\,(\boldsymbol{m}+e_{1}+e_{2}\,s)+(\boldsymbol{e}_{i-1,j}\,s)\,]_{Q}\end{split}$
where $\boldsymbol{e}_{i-1,j}\,s$ involves polynomial multiplication. Thus, the error spreads across all $N$ coefficients. Since the secret key $s$ has small coefficients (much less than $q_{0}$), each coefficient in the result incurs an error of magnitude proportional to $e_{j}$, but this error is now distributed. The maximum magnitude is bounded by $|e_{j}|\cdot\|s\|_{\infty}$.
III-D3 Bit flip in the Plaintext Before Decoding
If a bit flip is introduced in the coefficient $i$ of a plaintext $\boldsymbol{m}$, after decryption but before decoding, the effect is analogous to the case analyzed in Section III-C, since the plaintext being decoded is equivalent to:
$\boldsymbol{m}^{\prime}=\boldsymbol{m}+\boldsymbol{e}_{i-1,j}$ (15)
@@ III-E Theoretical Estimation of Error in RNS
$r_{k}=p\bmod q_{k}\quad\text{for }k=0,\dots,L-1$
When applied to all coefficients of a polynomial (either in the plaintext or in the ciphertext), this yields $L$ residue polynomials, commonly referred to as limbs. Reconstruction via the Chinese Remainder Theorem (CRT), as shown in Mohan [21, Chapter 5], is:
@@
where $Q_{k}=Q/q_{k}$ and $\frac{1}{Q_{k}}$ are the multiplicative inverse of $Q_{k}$.
A bit flip on the $j^{th}$ bit affecting the $k$-th residue changes it to $r_{k}^{\prime}=r_{k}+e_{k,j}$, which introduces an error into the reconstructed value:
$p^{\prime}-p=e_{k,j}\,Q_{k}\,\left[\left(\frac{1}{Q_{k}}\right)\bmod q_{k}\right]\pmod{Q}$ (17)
This error can be as large as $Q_{k}$, potentially spanning hundreds or thousands of bits, even though the bit flip originally affected a single bit in a single residue.
@@ III-F Theoretical Estimation of Error in NTT
To accelerate encoding and decoding operations, CKKS uses the negacyclic Number Theoretic Transform (NTT) over the ring $\mathbb{Z}_{q}[X]/(X^{N}+1)$. Let $\psi\in\mathbb{Z}_{q}$ be a primitive $2N$-th root of unity satisfying:
$\psi^{2}\equiv\xi,\quad\psi^{N}\equiv-1\pmod{q}$,
where $\xi$ is a primitive $N$-th root of unity. The NTT matrix is defined as:
We denote its inverse as $W^{-1}_{NTT}$. When a single bit flip occurs at bit position $j^{th}$ within coefficient $i$, the induced perturbation after NTT transformation becomes $\mathbf{m}^{\prime}-\mathbf{m}=W_{NTT}\,\mathbf{e}_{i,j}$. Therefore, the $L_{2}$ norm of the error is:
This behavior parallels that of the Discrete Fourier Transform (DFT), except that the NTT operates over $\mathbb{Z}_{q}$ with moduli $q$ typically of 306030-60 bits. The twiddle factors, being roots of unity modulo $q$, can amplify errors during the transformation, thereby increasing the error.
Figure 6 shows the values of $L_{2}(i,j)$ for $N=4$ and 64-bit coefficients in the NTT case. The horizontal axis indicates the bit-flip position $i\times 64+j$, and the vertical axis shows $L_{2}(i,j)$ on a logarithmic scale.
Figure 6: $L_{2}$-norm of the error after a single bit-flip in the plaintext during encoding. The horizontal axis shows the modified bit position $i\times 64+j$; the vertical axis shows the $L_{2}(i,j)$ norm on a logarithmic scale. Ring dimension: $N=4$.
As Figure 6 shows, even flips in the least significant bits cause the norm (error) to exceed four orders of magnitude–producing errors large enough to make correct decoding practically impossible, as will be further illustrated in Section V.
III-G Summary
Bit flip in the plaintext before encryption
the error remains localized in a single coefficient (the one where the bit flip occurred), with a magnitude proportional to $e_{j}=2^{j}$, depending on the position $j$ of the flipped bit.
Bit flip in the ciphertext ($c_{0}$)
identical to the plaintext case–the error remains localized after decryption, with magnitude proportional to the bit position $j$.
Bit flip in the ciphertext ($c_{1}$)
the error is multiplied by the secret key $s$, becoming distributed across all $N$ coefficients; its maximum magnitude is $\mathcal{O}(|e_{j}|\cdot\|s\|_{\infty})$.
Bit flip in the plaintext before decoding
behaves like a localized error in a single coefficient of the encoded polynomial–effectively equivalent to the encoding case, plus minor encryption noise.
RNS
a bit flip in a single limb (residue) can result in a reconstruction error of order $\mathcal{O}(Q_{k})$, where $Q_{k}=Q/q_{k}$ is typically a large value.
Negacyclic NTT
@@ III-G Summary
IV Methodology
@@
This section describes the methodology used to experimentally quantify the resilience of the CKKS scheme under single bit-flip fault injections.
IV-A Scheme Definition and Operational Assumptions
We consider four experimental configurations using CKKS:
  • Vanilla mode: All low-level performance optimizations are disabled; arithmetic occurs in a single large modulus in coefficient space (no RNS–single limb–and no NTT domain representation).
  • RNS-only mode: RNS decomposition is enabled (multi-limb representation), while NTT is bypassed.
  • NTT-only mode: NTT is enabled to accelerate polynomial multiplication / convolution, while RNS decomposition is disabled (single-limb modulus).
  • RNS+NTT mode: Fully optimized configuration (default in OpenFHE) with both RNS decomposition and NTT transform.
This setup allows us to isolate the individual and combined effects of RNS and NTT optimizations on fault propagation and resilience, while maintaining compatibility with standard security parameters.
In preliminary tests, we observed that both the NTT-only and the RNS+NTT configurations exhibit similar error amplification even for single-bit faults, in line with our theoretical analysis (Section III). In other words, a single bit flip at any stage of these two cases leads to undecryptable results. As a result, we focus our analysis on the other two cases (vanilla and RNS-only modes), where more interesting behavior can be observed and meaningfully characterized.
IV-B Fault / Error Injection Model
As in our theoretical analysis (Section III-A), we assume a transient single-bit fault model, motivated by soft errors such as radiation-induced bit flips in processor registers or memory buffers. Specifically, we inject a single bit flip in one coefficient of the output of a given operation during each execution. In the case of Encode and Decode, we inject the fault into the plaintext representation; for Encrypt and Decrypt, we inject it into the ciphertext. This approach models a fault affecting the destination register or output memory of these functions, without modifying the inputs or intermediate computations.
Figure 7: An input is first encoded and then encrypted. A single bit flip is applied to one coefficient at a time and at one bit position. The ciphertext is then decrypted and decoded, and the resulting output is compared against the unmodified reference output using the $L_{2}$-norm. For illustration, the ciphertext is represented as a single polynomial.
Figure 7 illustrates how the fault is injected at a single stage per experiment, directly on the polynomial representation (plaintext or ciphertext), simulating a post-operation corruption that propagates through subsequent stages of the computation.
IV-C Implementations and Instrumentation
IV-C1 CKKS Libraries
We use two different CKKS implementations: HEAAN 1.0. First publicly available CKKS implementation. This specific version does not support RNS and NTT, which makes it suitable for the evaluation of the vanilla mode. It uses arbitrary-precision coefficients managed by the NTL library. OpenFHE 1.3.0: highly-optimized open-source (de facto) standard implementation of CKKS [22]. It natively supports RNS and NTT, using 64-bit internal coefficients.
@@ IV-C1 CKKS Libraries
We minimally adapt both libraries to gain complete control over the selection of pseudo-random number generator (PRNG) seeds. This ensures strict reproducibility of experiments and allows precise isolation of the effects caused by individual bit flips.
In the case of HEAAN 1.0, we consider the following implementation details:
@@
  • The bit length of each plaintext coefficient is not fixed in advance; it depends on internal combinations of $q_{0}$ and $\Delta$, and may vary depending on the PRNG seed used by the scheme.
  • The ciphertext bit length is of length $q_{0}$.
@@
  • Due to NTL implementation constraints, bit positions larger than the $Q$ of the coefficient are ignored in internal operations, which limits the scope of bit-level analysis.
In the case of OpenFHE, we consider the following implementation details:
  • Use of 64-bit integer coefficients (uint64_t) without arbitrary-precision arithmetic.
@@
  • The client can configure the initial modulus $q_{0}$. This parameter determines the starting bit-level precision of the scheme. The subsequent moduli $q_{i}$ are automatically selected to match the bit length of $q_{0}$, ensuring uniform precision across levels.
  • The plaintext and the ciphertext belong to the same ring. They both have polynomial coefficients of size $q_{0}$, regardless of the scaling factor $\Delta$.
@@
  • The total number of moduli depends on the required multiplicative depth of the computation. One moduli for each RNS limb.
@@
  • 128-bit arithmetic support, which–though not used in this work–offers a pathway for future experiments.
IV-C2 Using OpenFHE as Vanilla CKKS Implementation
OpenFHE is substantially more computationally efficient than HEAAN. Consequently, we primarily use OpenFHE in this work, although selected experiments are replicated in HEAAN for validation purposes. However, it does not natively offer a mechanism to disable RNS and NTT optimizations to enable OpenFHE on vanilla mode. We explicitly bypass both RNS and NTT optimizations. RNS: We enforced a single-limb representation, which is feasible for depth-zero operations (i.e., no ciphertext multiplications). Additionally, automatic rescaling was disabled and replaced with manual scaling to prevent the introduction of additional limbs. NTT: Before injecting a bit flip, we move the polynomial from evaluation space back to coefficient space using the inverse NTT (INTT). We flip the desired bit and transform the polynomial back to the evaluation space using NTT. After this hook, the pipeline continue normally.
@@ IV-C2 Using OpenFHE as Vanilla CKKS Implementation
Since the NTT operates modulo $q_{0}$, flipping bits beyond the bit width of $q_{0}$ causes modular wraparound, which can alter the intended effect of the bit flip and potentially lead to catastrophic decryption errors. To minimize distortion in the analysis, we set $q_{0}$ to the maximum value supported by the library–60 bits–ensuring that most bit flips occur within the valid range and reducing unintended side effects from modular arithmetic.
IV-C3 Validation of Vanilla Mode using OpenFHE
To ensure that vanilla OpenFHE mode faithfully replicates the behavior of HEAAN 1.0, we follow these validation steps, 1) Configure OpenFHE on vanilla mode (disabling RNS and NTT) and set $q_{0}$ and $\Delta$ to match HEAAN’s effective bit precision, 2) Restrict input values in HEAAN to real numbers, in line with OpenFHE’s input constraints, 3) Compare $L_{2}$-norms and error statistics after identical bit flips are introduced in both libraries and 4) Verify consistent sensitivity patterns across low- and high-order bit positions.
IV-D Experimental Procedure
In OpenFHE, ciphertexts and plaintexts are internally represented using NTT transformation (evaluation domain). Since we need to inject faults in the coefficients of the polynomials, we must ensure that we are working in the coefficient domain. For the vanilla and RNS modes, we first apply the inverse NTT to bring the polynomial back to the coefficient domain, then we perform the bit flip, and finally we reapply the NTT to restore the polynomial to the evaluation domain.
For each bit flip experiment, we follow these steps:
@@ IV-D Experimental Procedure
    Apply the inverse NTT (INTT) to bring the polynomial into coefficient space.
@@
    Reapply NTT to return the polynomial to evaluation space.
@@
We repeat this process for every bit position in each coefficient of the different stages of the CKKS pipeline.
Input Vector Generation and PRNG Control
Our analysis is application-agnostic, employing input vectors of random real numbers drawn from a uniform distribution. To assess the sensitivity of the CKKS scheme, we sample inputs from various intervals whose boundaries are defined by powers of two. Unless otherwise specified, input values are sampled as uniformly distributed real numbers within the interval 0 to 2560256, and represented in double-precision floating-point format. The bounded input range enables testing multiple scaling factors ($\Delta$), allowing for a more comprehensive analysis of their impact on encoding and rounding errors. We analyze the trade-off between precision and noise growth by varying $\Delta$ across a wide range, ensuring it stays within the coefficient bounds detailed in Section II-B.
To evaluate stability impact of the seed variation, we repeat each experiment (i.e., each bit flip) 2500 times. Each repetition involves 100 different seeds for the scheme’s PRNG and 25 for the input generator. These seed combinations are fixed and shared across all experiments, ensuring a fair comparison using equivalent input conditions. The variability observed across repetitions is negligible, with no discernible differences in the resulting patterns, indicating that the adopted sampling strategy is sufficient for the purposes of this study.
Number of Coefficients
In order to ensure reasonable computation times given the number of seeds, we first studied the effect of the ring dimension $N$ on bit-flip sensitivity patterns. This parameter has the greatest impact on computational performance: a larger $N$ provides better security, but at the cost of increased computation. Since the bit-flip sensitivity patterns were found to be consistent across experiments with varying ring dimensions $N$ (from $2^{3}$ up to $2^{16}$), we restrict our analysis to smaller values of $N$. This choice both improves performance and facilitates visualization: it is far easier to discern structure in the behavior of a small set of coefficients in a single plot than to interpret patterns spread across thousands simultaneously.
Analyzed Stages
As presented in Section III, we corroborate experimentally that flipping a single bit in the plaintext during the Encode or Decode stages has the same effect as flipping a bit in the first polynomial ($c_{0}$) of the ciphertext during Encrypt or Decrypt. For this reason, we focus our study on the behavior of the ciphertext under bit-flip faults. This approach allows us, within a single experiment, to analyze both the behavior of the first polynomial (and thus the plaintext) and the behavior of the second polynomial of the ciphertext.
IV-E Evaluation Metrics
Given the original decrypted vector $\mathbf{x}$ (without bit flips) and the resulting vector $\mathbf{y}$ after error injection and decryption, we quantify the deviation introduced by each bit flip by two complementary evaluation metrics: